OpenStudy (anonymous):

A concert hall earned $40,000 in sold-out comedy ticket sales. Front section tickets cost $150 per seat and back section tickets cost $50 per seat. The number of front section seats is three times the number of back section seats. How many seats are in the front section?

OpenStudy (anonymous):

150f + 50b = 40000 f = = 3b => 150f + 50f/3 = 40,000 => 500f/3 = 40,000 => f = 120000/500 => f = 240

OpenStudy (anonymous):

I did solve this problem before !